15. DO NOT DO THE GRAPH OR THE DOMAIN OR RANGE PART. THOSE DO NOT COUNT....

15.

DO NOT DO THE GRAPH OR THE DOMAIN OR RANGE PART. THOSE DO NOT COUNT. WE ARE NOT DOING THEM IN MATH 1003.Find the factors that are common in the numerator and the denominator. Then find the intercepts and asymptotes. (If an answer does not exist, enter DNE. Enter your asymptotes as a comma-separated list of equations if necessary.)r(x) =

x2 + 8x − 9
x2 + 3x − 4
x-intercept    

(x, y)

=
  
y-intercept    

(x, y)

=
  
vertical asymptote(s)    
horizontal asymptote    
Answer #1

aty x² +8x-9 (xt) Corto) x² + 3x-4 (26)(x+h) - xétencept put y=0 0-9,0) Y intercept put x = 0 (0,9/4) asymptotes veutical at